# Gamma-ray burst collimated fireballs: long vs short bursts by angular momentum and mass ratio

### v2

@@ 7.2 Collimated fireball angle $\theta_{\rm coll}$ relates to jet angle
As depicted in Figs. 4 and 5, the collimated fireball of photon-pair plasma energy density $\rho_{\gamma}$ shows denser and wider at the base $(\theta\rightarrow\pi)$, while less dense and narrower at the top $(\theta\rightarrow 0)$. It is the consequence of gravitational binding potential $\phi$ and centrifugal potential $\phi_{c}$ in the temperature profile $T_{\gamma}$ (6.5) at the end of homologous collapses. The opaque collimated fireball pressure (density) $p_{\gamma}\sim\rho_{\gamma}\propto T_{\gamma}^{4}$ is no longer isotropic and its maximal value along $\hat{z}$-axis, decreasing from $R=0$ to $R=R_{\gamma}$. In the transverse directions perpendicular to $\hat{z}$, its value decreases to the collimated fireball boundary (orange thick line in Fig. 5). Namely, most collimated fireball energy power concentrates at the centre along $\hat{z}$.
@@
On the contrary, the collimated fireball’s angular momentum $j\rho_{\gamma}dV$ vanishes along $\hat{z}$ direction ($\theta=0$) from $R=0$ to $R_{\gamma}$, and increases in the transverse directions ($\theta=\pi/2$). Namely, most fireball angular momentum is carried near the collimated fireball boundary. An outward thermal pressure in the fireball centre along the $\hat{z}$ direction and the conservation of angular momentum most near the collimated fireball boundary possibly leads to an ultra-relativistic jet.
@@
In the phase of early jet launching, we argue that such a collimated fireball will self-similarly evolve to the normal jet profile of energy density denser and wider at the top $(\theta\rightarrow 0)$, less dense and narrower at the base $(\theta\rightarrow\pi)$. The reason is that the highly dense and energetic part is deeply opaque and has a large thermal pressure $p_{\gamma}$ along the $\hat{z}$ direction, ultra-relativistically moving outward. As a result, denser and more energetic parts pile up on the front, moving at almost light speed, while dilute and less energetic parts leave behind as a narrow tail. It is also the case for the collimated fireball outflow engulfing a small amount of material. This situation is analogous to the spherical case of ultra-relativistically hydrodynamic expansion: a spherical fireball evolves to a self-similar (homologous) spherical slab at the front moving at almost light speed, and a tail of dilute number and energy densities left behind, which are studied both analytically and numerically [18, 19].
@@
We must mention the condition of a collimated fireball forming an outward propagating jet. Figure 5 shows collimated fireballs form inside external baryon shells (green layers). Whether or not they can erupt out of external baryon shells requires the ratio $E_{\gamma}/M_{\rm ej}\gg 1$, see the references and discussions for Eq. (8.4), where $E_{\gamma}$ is the total energy of collimated fireballs and $M_{\rm ej}$ is the ejecta mass contributing from the external baryon shell. The $M_{\rm ej}$ is much smaller than the total baryon core mass $M$, i.e., $M_{\rm ej}\ll M$, because Figure 1 shows the baryon external shell mass density is very small for $r>10$ comparing the core centre density. While collimated fireball energies are about $E_{\gamma}/M\sim 10^{-2}$ (6.16). This implies the possibility of collimated fireballs erupting from baryon cores and forming ultra-relativistic ejecta with nontrivial angular momenta.
@@
where $R^{j}_{\gamma}(\theta=\pi/2)$ is the maximal lateral size of collimated fireballs and $R^{j}_{\gamma}(\theta=0)$ is the corresponding radius of the spherical fireball. Such an angle $\theta_{\rm coll}$ characterizes the axial symmetric collimation of photon-pair fireball along the $\hat{z}$ direction before its ultra-relativistic hydrodynamic outflow.
@@ 8 Collimated fireball intrinsic correlations
$(J/M,\rho^{\rm in}_{c}/\rho_{0},\gamma,\lambda),$ (8.1)
where the total mass and angular momentum ratio $J/M$ characterizes rotating systems in gravitational collapsing processes, the initial centre density $\rho^{\rm in}_{c}/\rho_{0}$ crucially relates to the gravitational binding energy gain, averaged thermal index $\gamma$ and material binding energy $\lambda$ represent material reactions in processes. The photon-pair collimated fireball energetic properties (6.8,6.9) and (6.13,6.15) explicitly depend on the initial and final core central densities $\rho_{c}^{\rm in}$ and $\rho_{c}^{\rm fi}=\rho_{c}^{\rm max}\approx 10\rho_{0}$. They determine the gravitational binding energy gain and conversion to baryon kinetic energy and photon-pair thermal energy in gravitational collapses. These characteristics should also be held for gravitational collapses of non-spherical cores or binary mergers by introducing effective initial and final central densities $(\rho_{c}^{\rm in})_{\rm eff}$ and $(\rho_{c}^{\rm max})_{\rm eff}$ of the system.
@@ 8.1 Intrinsic correlations among collimated fireball quantities
All collimated fireball quantities (6.8,6.9) and (6.13,6.14,6.15) are expressed as scaling laws of the initial core centre density $\rho_{c}^{\rm in}$ since it characterizes how much the gravitational binding energy gains when the systems of masses $M$ gravitationally collapse from $\rho_{c}^{\rm in}$ to final core centre density $\rho_{c}^{\rm fi}=\rho_{c}^{\rm max}\approx 10\rho_{0}$.
By simplifying out the initial core centre density $\rho_{c}^{\rm in}$ among Eqs. (6.8,6.9) and (6.13,6.15), we obtained theoretical correlations among the spectral peak $E_{p}$, total photon-pair energy $E_{\gamma}$ and peak luminosity $L_{\gamma}\propto cR_{\gamma}^{2}\langle T_{\gamma}\rangle^{4}\propto cE_{\gamma}/R_{\gamma}$ of the photon-pair sphere or collimation in a local rest frame,
@@
depending on the unique parameter $\gamma\gtrsim 1$ and $\chi=\frac{\gamma-1}{4\gamma-5}\sim{\mathcal{O}}(1)$  [^4] It is an inverse of the original $\chi-$definition [26]..
@@
We cannot analytically obtain intrinsic correlations between the collimated fireball angle $\theta_{\rm coll}$ (7.2), the ratio $J/M$ and other collimated fireball quantities $E_{p}^{j}$, $E_{\gamma}^{j}$, $L_{\gamma}^{j}$ and $R_{\gamma}^{j}$. However, in the next Sec. 9, we will try to obtain the numerical trends (Table 1) of collimated fireball quantities in terms of rotation parameter $a^{2}\Omega^{2}_{c}\propto J/M$ at the end of homologous collapses. From these numerical trends, we infer collimated fireball intrinsic correlations, including the collimated fireball angle $\theta_{\rm coll}$. Henceforth, we omit the superscript j of collimated fireball quantities for simplifying notations unless otherwise specified.
@@ 8.2 Intrinsic correlations imprint on observed GRB data
The theoretically obtained intrinsic correlations (8.6) and (8.7) seem to give the Amati and Yonetoku empirical correlations in GRBs [45, 46], if we identify the photon-pair collimated fireball energy $E_{\gamma}$ and mean energy $\langle T_{\gamma}\rangle$ to the observed isotropic equivalent energy $E_{\rm iso}$ and peak energy $E_{p}$. However, we cannot simply make such identification for the reasons that a photon-pair collimated fireball is opaque and undergoes complex processes up to the transparency at which it becomes an observably relevant photosphere or collimated one. One of the reasons is that the peak energy $E_{p}$ value sensitively depends on the baryonic material loading (ejecta mass) $M_{\rm ej}$ on collimated fireballs, see for example, Refs. [18, 19, 20, 21]. However, we present a general discussion on the connections between theoretical correlations (8.6-8.11) and trends (Table 1) of different collimated fireballs and observational correlations of different GRB sources.
@@
$0.001$ $13.2^{\circ}$ $1.01$ $0.998$ $4.29$ $6.9/30$ $0.99$ $0.92$
$0.005$ $7.8^{\circ}$ $1.25$ $0.99$ $7.24$ $4.1/30$ $0.97$ $0.63$
$0.01$ $6.3^{\circ}$ $1.38$ $0.98$ $8.90$ $3.3/30$ $0.96$ $0.52$
$\Rightarrow 0.02$ $4.8^{\circ}$ $1.52$ $0.97$ $11.6$ $2.5/30$ $0.93$ $0.42$
$0.03$ $4.2^{\circ}$ $1.60$ $0.95$ $12.9$ $2.2/30$ $0.91$ $0.37$
$0.04$ $3.8^{\circ}$ $1.66$ $0.94$ $14.1$ $2.0/30$ $0.89$ $0.33$
$0.05$ $3.4^{\circ}$ $1.71$ $0.93$ $15.5$ $1.8/30$ $0.88$ $0.31$
$0.06$ $3.2^{\circ}$ $1.74$ $0.92$ $16.2$ $1.7/30$ $0.86$ $0.29$
$0.07$ $3.1^{\circ}$ $1.78$ $0.91$ $17.1$ $1.6/30$ $0.85$ $0.27$
$0.08$ $2.9^{\circ}$ $1.81$ $0.90$ $18.0$ $1.5/30$ $0.84$ $0.26$
$0.09$ $2.7^{\circ}$ $1.84$ $0.89$ $19.1$ $1.4/30$ $0.83$ $0.25$
$0.10$ $2.5^{\circ}$ $1.86$ $0.88$ $20.3$ $1.3/30$ $0.82$ $0.24$
$0.15$ $2.1^{\circ}$ $1.95$ $0.85$ $23.2$ $1.1/30$ $0.78$ $0.20$
Table 1: For selected values of the rotation parameter $a^{2}\Omega^{2}_{c}$, which is proportionally related to the ratio $J/M$ (5.30), we tabulate the collimated angle $\theta_{\rm coll}$ (7.2) and the ratios of the spectral peak $E^{j}_{p}/E_{p}$, total energy $E^{j}_{\gamma}/E_{\gamma}$, maximal luminosity $L^{j}_{\gamma}/L_{\gamma}\propto(E^{j}_{\gamma}/E_{\gamma})(R_{\gamma}/R^{j}_{\gamma})$, minimal size $R^{j}_{\gamma}/R_{\gamma}$, total particle number $N^{j}_{\gamma}/N_{\gamma}$ and volume $V^{j}_{\gamma}/V_{\gamma}$ between the photon-pair spherical and collimated fireballs for comparisons. The collimated (with superscript $j$) and spherical fireball quantities are given by Eqs. (6.8,6.14,6.15). Their ratios do not explicitly depend on the initial (final) core centre density $\rho_{c}^{\rm in}$ ($\rho_{c}^{\rm fi}$) of gravitational collapses.

### v3

@@ 7.2 Collimated fireball angle $\theta_{\rm coll}$ relates to jet angle
As depicted in Figs. 4 and 5, the collimated fireball of photon-pair plasma energy density $\rho_{\gamma}$ shows denser and wider at the base $(\theta\rightarrow\pi)$, while less dense and narrower at the top $(\theta\rightarrow 0)$. It is the consequence of the gravitational binding potential $\phi$ and centrifugal potential $\phi_{c}$ in the temperature profile $T_{\gamma}$ (6.5) at the end of homologous collapses. The opaque collimated fireball pressure (density) $p_{\gamma}\sim\rho_{\gamma}\propto T_{\gamma}^{4}$ is no longer isotropic and its maximal value along $\hat{z}$-axis, decreasing from $R=0$ to $R=R_{\gamma}$. In the transverse directions perpendicular to $\hat{z}$, its value decreases to the collimated fireball boundary (orange thick line in Fig. 5). Namely, most collimated fireball energy concentrates at the centre along $\hat{z}$.
@@
On the contrary, the collimated fireball’s angular momentum $j\rho_{\gamma}dV$ vanishes along $\hat{z}$ direction ($\theta=0$) from $R=0$ to $R_{\gamma}$, and increases in the transverse directions ($\theta=\pi/2$). Namely, most fireballs’ angular momentum is carried near the collimated fireball boundary. An outward thermal pressure in the fireball centre along the $\hat{z}$ direction and the conservation of angular momentum near the collimated fireball boundary possibly leads to an ultra-relativistic jet.
@@
In the phase of early jet launching, we argue that such a collimated fireball will self-similarly evolve to the normal jet profile of energy density denser and wider at the top $(\theta\rightarrow 0)$, less dense and narrower at the base $(\theta\rightarrow\pi)$. The reason is that the highly dense and energetic part is deeply opaque and has a large thermal pressure $p_{\gamma}$ along the $\hat{z}$ direction, ultra-relativistically moving outward. As a result, denser and more energetic parts pile up on the front, moving at almost light speed, while dilute and less energetic parts leave behind a narrow tail. It is also the case for the collimated fireball outflow engulfing a small amount of material. This situation is analogous to the spherical case of ultra-relativistically hydrodynamic expansion: a spherical fireball evolves to a self-similar (homologous) spherical slab at the front moving at almost light speed, and a tail of dilute number and energy densities left behind, which are studied both analytically and numerically [18, 19].
@@
We must mention the condition of a collimated fireball forming an outward-propagating jet. Figure 5 shows that collimated fireballs form inside external baryon shells (green layers). Whether or not they can erupt out of external baryon shells requires the ratio $E_{\gamma}/M_{\rm ej}\gg 1$, see the references and discussions for Eq. (8.4), where $E_{\gamma}$ is the total energy of collimated fireballs and $M_{\rm ej}$ is the ejecta mass contributing from the external baryon shell. The $M_{\rm ej}$ is much smaller than the total baryon core mass $M$, i.e., $M_{\rm ej}\ll M$, because Figure 1 shows the baryon external shell mass density is very small for $r>10$ comparing the core centre density. While collimated fireball energies are about $E_{\gamma}/M\sim 10^{-2}$ (6.17). This implies the possibility of collimated fireballs erupting from baryon cores and forming ultra-relativistic ejecta with nontrivial angular momenta.
@@
where $R^{j}_{\gamma}(\theta=\pi/2)$ is the maximal lateral size of collimated fireballs and $R^{j}_{\gamma}(\theta=0)$ is the corresponding radius of the spherical fireball. Such an angle $\theta_{\rm coll}$ characterizes the axially symmetric collimation of photon-pair fireball along the $\hat{z}$ direction before its ultra-relativistic hydrodynamic outflow.
@@
We make some remarks to end this section. The collimated fireball energy $E_{\gamma}$ and ejector mass $M_{\rm ej}$ ratio $\Gamma_{0}\approx E_{\gamma}/M_{\rm ej}>1$ (8.3) is one of the necessary conditions for a collimated fireball erupting out of external baryon shells (stellar envelope) as an outward collimated flow, accompanied by cocoon configurations [60]. However, to maintain a distinct “thin” jet structure in its ultra-relativistic propagation and penetrate through the stellar envelope, the collimated outflow should be energetic $E_{\gamma}/M_{\rm ej}\gg 1$ and carry large enough angular momentum $j_{\gamma}/M$ (7.1) to overcome small shearing forces (torques) from the surrounding material. It is the case particularly for large $M$ and $J/M$ binary coalescence (short GRBs), producing small $\theta_{\rm coll}$ “thin” jets, which carry away more angular momentum along the rotation axes, and interact with less massive $M_{\rm ej}$ stellar envelopes, resulting in weak cocoons, more disc-like configurations and strong gravitational wave signals.
Otherwise, for small values $E_{\gamma}/M_{\rm ej}>1$ and $j_{\gamma}/M$, outward collimated flows become relativistic shocked jets, due to dissipation of their energy and angular momentum to the surrounding material (Newtonian shocked stellar material), creating cocoon configurations. It is the case particularly for massive $M$ and small $J/M$ core collapses (long GRBs), producing large $\theta_{\rm coll}$ “fat” jets interacting with more massive $M_{\rm ej}$ stellar envelopes, resulting in strong cocoons, less disc-like configurations and weak gravitational wave signals.
The fractions of shocked jet and stellar material components in cocoons depend on how collimated fireballs transfer their energies $E_{\gamma}$ and angular momenta $j_{\gamma}$ to the stellar envelope materials. The shocked jet angle $\theta_{\rm jet}$ should be smaller than the collimated fireball angle $\theta_{\rm coll}$, because of angular momentum transfers. Less energetic collimated fireballs $E_{\gamma}/M_{\rm ej}\gtrsim 1$ become “failed” jets, creating hot spots of Newtonian shocked stellar materials. These cases should be the most probable fates of massive stellar core collapses, but their low-frequency signals are probably too faint for recent observations.
@@ 8 Collimated fireball intrinsic correlations
$(J/M,\rho^{\rm in}_{c}/\rho_{0},\rho^{\rm fi}_{c}/\rho_{0},\gamma,\lambda),$ (8.1)
where the total mass and angular momentum ratio $J/M$ characterizes rotating systems in gravitational collapsing processes, the initial and final centre densities $\rho^{\rm in}_{c}/\rho_{0}$ and $\rho^{\rm fi}_{c}/\rho_{0}$ crucially relate to the gravitational binding energy gain, averaged thermal index $\gamma$ and material binding energy $\lambda$ represent material reactions in processes. The photon-pair collimated fireball energetic properties (6.9,6.10) and (6.14,6.16) explicitly depend on the initial and final core central densities $\rho_{c}^{\rm in}$ and $\rho_{c}^{\rm fi}=\rho_{c}^{\rm max}\approx 10\rho_{0}$. They determine the gravitational binding energy gain and conversion to baryon kinetic energy and photon-pair thermal energy in gravitational collapses. These characteristics should also be held for gravitational collapses of non-spherical cores or binary mergers by introducing effective initial and final central densities $(\rho_{c}^{\rm in})_{\rm eff}$ and $(\rho_{c}^{\rm max})_{\rm eff}$ of the system. For appropriate configurations (8.1), this scenario could also provide an efficient mechanism of photon-pair driven supernova explosion, and detailed investigations are required.
@@ 8.1 Intrinsic correlations among collimated fireball quantities
All collimated fireball quantities (6.9,6.10) and (6.14,6.15,6.16) are expressed as scaling laws of the initial core centre density $\rho_{c}^{\rm in}$ since it characterizes how much the gravitational binding energy gains when the systems of masses $M$ gravitationally collapse from $\rho_{c}^{\rm in}$ to final core centre density $\rho_{c}^{\rm fi}=\rho_{c}^{\rm max}\approx 10\rho_{0}$.
By simplifying out the initial core centre density $\rho_{c}^{\rm in}$ among Eqs. (6.9,6.10) and (6.14,6.16), we obtained theoretical correlations among the spectral peak $E_{p}$, total photon-pair energy $E_{\gamma}$ and peak luminosity $L_{\gamma}\propto cR_{\gamma}^{2}\langle T_{\gamma}\rangle^{4}\propto cE_{\gamma}/R_{\gamma}$ of the photon-pair sphere or collimation in a local rest frame,
@@
which depend on the unique parameter $\gamma\gtrsim 1$ and $\chi=\frac{\gamma-1}{4\gamma-5}\sim{\mathcal{O}}(1)$  [^4] It is an inverse of the original $\chi-$definition [36]..
@@
We cannot analytically obtain intrinsic correlations between the collimated fireball angle $\theta_{\rm coll}$ (7.2), the ratio $J/M$ and other collimated fireball quantities $E_{p}^{j}$, $E_{\gamma}^{j}$, $L_{\gamma}^{j}$ and $R_{\gamma}^{j}$. However, in the next Sec. 9, we will try to obtain the numerical trends (Table 1) of collimated fireball quantities in terms of rotation parameter $a^{2}\Omega^{2}_{c}\propto J/M$ at the end of homologous collapses. From these numerical trends, we infer collimated fireball intrinsic correlations, including the collimated fireball angle $\theta_{\rm coll}$. We omit the superscript j of collimated fireball quantities for simplifying notations unless otherwise specified.
@@ 8.2 Intrinsic correlations imprint on observed GRB data
The theoretically obtained intrinsic correlations (8.6) and (8.7) seem to give the Amati and Yonetoku empirical correlations in GRBs [62, 63], if we identify the photon-pair collimated fireball energy $E_{\gamma}$ and mean energy $\langle T_{\gamma}\rangle$ to the observed isotropic equivalent energy $E_{\rm iso}$ and peak energy $E_{p}$. However, we cannot simply make such an identification for the reasons that a photon-pair collimated fireball is opaque and undergoes complex processes up to the transparency at which it becomes an observably relevant photosphere or collimated one. One of the reasons is that the peak energy $E_{p}$ value sensitively depends on the baryonic material loading (ejecta mass) $M_{\rm ej}$ on collimated fireballs, see for example, Refs. [18, 19, 20, 21]. However, we present a general discussion on the connections between theoretical correlations (8.6-8.11) and trends (Table 1) of different collimated fireballs and observational correlations of different GRB sources.
@@
We end this section by emphasizing that the GRB physics is most complex as various dynamic processes are involved in time evolution. The central engine physics at the initial explosion could be simpler, but masked by surrounding opaque material, cannot be directly probed by observations. In theoretical correlations (8.6-8.11), the universal scaling laws should be primary, and the relations between scaling indexes can receive corrections. Therefore, the verifications of theoretical correlations (8.6-8.11) and correlations relating to jet angles discussed below must be conducted carefully with caution. They cannot be justified or falsified by a limited number of observations. We like to mention recent empirical studies [64, 65] of GRBs’ luminosity and duration correlation, similar to Eq. (8.11) or Eq. (7.7) of Ref. [36].
@@
0.0010.0010.001 $13.2^{\circ}$ 1.011.011.01 0.9980.9980.998 4.294.294.29 6.9/306.9/306.9 / 30 0.990.990.99 0.920.920.92
0.0050.0050.005 $7.8^{\circ}$ 1.251.251.25 0.990.990.99 7.247.247.24 4.1/304.1/304.1 / 30 0.970.970.97 0.630.630.63
0.010.010.01 $6.3^{\circ}$ 1.381.381.38 0.980.980.98 8.908.908.90 3.3/303.3/303.3 / 30 0.960.960.96 0.520.520.52
0.02$\Rightarrow$ 0.02⇒ 0.02 $4.8^{\circ}$ 1.521.521.52 0.970.970.97 11.611.611.6 2.5/302.5/302.5 / 30 0.930.930.93 0.420.420.42
0.030.030.03 $4.2^{\circ}$ 1.601.601.60 0.950.950.95 12.912.912.9 2.2/302.2/302.2 / 30 0.910.910.91 0.370.370.37
0.040.040.04 $3.8^{\circ}$ 1.661.661.66 0.940.940.94 14.114.114.1 2.0/302.0/302.0 / 30 0.890.890.89 0.330.330.33
0.050.050.05 $3.4^{\circ}$ 1.711.711.71 0.930.930.93 15.515.515.5 1.8/301.8/301.8 / 30 0.880.880.88 0.310.310.31
0.060.060.06 $3.2^{\circ}$ 1.741.741.74 0.920.920.92 16.216.216.2 1.7/301.7/301.7 / 30 0.860.860.86 0.290.290.29
0.070.070.07 $3.1^{\circ}$ 1.781.781.78 0.910.910.91 17.117.117.1 1.6/301.6/301.6 / 30 0.850.850.85 0.270.270.27
0.080.080.08 $2.9^{\circ}$ 1.811.811.81 0.900.900.90 18.018.018.0 1.5/301.5/301.5 / 30 0.840.840.84 0.260.260.26
0.090.090.09 $2.7^{\circ}$ 1.841.841.84 0.890.890.89 19.119.119.1 1.4/301.4/301.4 / 30 0.830.830.83 0.250.250.25
0.100.100.10 $2.5^{\circ}$ 1.861.861.86 0.880.880.88 20.320.320.3 1.3/301.3/301.3 / 30 0.820.820.82 0.240.240.24
0.150.150.15 $2.1^{\circ}$ 1.951.951.95 0.850.850.85 23.223.223.2 1.1/301.1/301.1 / 30 0.780.780.78 0.200.200.20
Table 1: For selected values of the rotation parameter $a^{2}\Omega^{2}_{c}$, which is proportionally related to the ratio $J/M$ (5.30), we tabulate the collimated angle $\theta_{\rm coll}$ (7.2) and the ratios of the spectral peak $E^{j}_{p}/E_{p}$, total energy $E^{j}_{\gamma}/E_{\gamma}$, maximal luminosity $L^{j}_{\gamma}/L_{\gamma}\propto(E^{j}_{\gamma}/E_{\gamma})(R_{\gamma}/R^{j}_{\gamma})$, minimal size $R^{j}_{\gamma}/R_{\gamma}$, total particle number $N^{j}_{\gamma}/N_{\gamma}$ and volume $V^{j}_{\gamma}/V_{\gamma}$ between the photon-pair spherical and collimated fireballs for comparisons. The collimated (with superscript $j$) and spherical fireball quantities are given by Eqs. (6.9,6.15,6.16). Their ratios do not explicitly depend on the initial (final) core centre density $\rho_{c}^{\rm in}$ ($\rho_{c}^{\rm fi}$) of gravitational collapses.